OpenStudy (anonymous):

find initial value problem (ivp) 2y=5x+dy/dx y(0)=1

OpenStudy (anonymous):

OK, so we're first going to put it in the standard form of y'+p(x)y=q(x): \[y'-2y=-5x\]We find an integrating factor, which will be \[u=e^{\int p(x)dx}=e^{-2x}\]and multiply both sides of the equation by it: \begin{align} u(y'-2y&)=u(-5x)\\ y' e^{-2x}-2ye^{-2x} &= -5xe^{-2x} \\ \left(ye^{-2x}\right)'&=-5xe^{-2x}\\ ye^{-2x}&=-5\int x e^{-2x}dx \\ ye^{-2x}&=-5\left(-\frac{1}{2}xe^{-2x}+\frac{1}{2}\int e^{-2x}dx\right)\\ ye^{-2x}&=-5\left(-\frac{1}{2}xe^{-2x}-\frac{1}{4}e^{-2x}+C\right)\\ ye^{-2x}&=\frac{5}{2}xe^{-2x}+\frac{5}{4}e^{-2x}+C\\ y&=\frac{5}{2}x+\frac{5}{4}+Ce^{2x} \end{align} Now we use the initial condition of y(0)=1 to solve for C: \[1=\frac{5}{4}+Ce^{2\cdot 0}=\frac{5}{4}+C\] Therefore C=-1/4 and your specific solution is \[y=\frac{5}{2}x+\frac{5}{4}-\frac{1}{4} e^{-2x}.\]